OpenStudy (anonymous):

What is the value of f(16) − f(0) when f(x) = 4x – 8?

OpenStudy (mathstudent55):

Find f(16) by replacing 16 for x in 4x - 8 Then do the same for 0. Then subtract the results.

OpenStudy (anonymous):

so it would be 56-8 which gives you the answer of 48

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

You are on the right track, but you need to be careful. \(f(x) = 4x - 8\) \(f(16) = 4(16) - 8 = 64 - 8 = 56\) \(f(0) = 4(0) - 8 = ~~?\) Then \(f(16) - f(0) = 56 - ~~? = ~~ ?\)

OpenStudy (mathstudent55):

Look above. What is f(0) equal to?

OpenStudy (anonymous):

0

OpenStudy (mathstudent55):

No. f(0) = 4(0) - 8 = 0 - 8 What is 0 - 8 equal to?

OpenStudy (anonymous):

oh sorry its equal to -8

OpenStudy (mathstudent55):

Right. Now remember you are subtracting the values of the function, so what is 56 - (-8) = ?

OpenStudy (anonymous):

64

OpenStudy (mathstudent55):

Correct.

OpenStudy (mathstudent55):

That's what I meant by being careful. You needed to do 56 - (-8) = 56 + 8 = 64, which you did now, not 56 - 8 = 48, which you did earlier.

OpenStudy (anonymous):

oh got it thank you so much

OpenStudy (mathstudent55):

You're welcome.